OpenStudy (anonymous):

The Smiths plan to recarpet their family room, which measures 15 ft. by 20 ft. How many square yards of carpet are needed?

OpenStudy (anonymous):

5 yds * 6 2/3 yds = 30 + 3 1/3 = 33 1/3 sq yds